OpenStudy (anonymous):

if someone has grades of 73 and 79 on two test and she wants an average of at least 71 what is the possible score. Less than 61 61 or less greater than 61 61 or greater.. I came up with 73+79+ x divided by 3 which gives me 152 so do I subtract from there

OpenStudy (anonymous):

The answer is 61. if you add 61+73+79 it equals 213 and then you divide by 3 and you get 71

OpenStudy (anonymous):

Does this make sense?

OpenStudy (anonymous):

So I would add all three. So she would need to have a 61 or greater?

OpenStudy (anonymous):

Right because if you added 60 instead her average would be 70.666.%. and if you added 62 to the other two numbers you would get 71.333 %. I hope this helps :)

OpenStudy (anonymous):

so to find any average you add up all the numbers and then divide by however many numbers you added together :) in this case it was 3 numbers you added together so thats why you divide by 3 after adding them up.

OpenStudy (texaschic101):

(73 + 79 + x) / 3 >= 71 (greater then or equal ) (152 + x) / 3 >= 71 -- multiply both sides by 3 152 + x >= 213 -- subtract 152 from both sides x >= 213 - 152 x >= 61........answer is 61 or greater